OpenStudy (anonymous):

what is 2 and 1/5+1 and 1/10? Will give medal!!!please help A)3 and 1/5 B)3 and 3/5 C)3 and 1/10 D)3 and 3/10

OpenStudy (kc_kennylau):

\[2\frac15+1\frac1{10}\\=2+\frac15+1+\frac1{10}\\=2+1+\frac15+\frac1{10}\\=3+\frac2{10}+\frac1{10}\\=3+\frac3{10}\\=3\frac3{10}\]

OpenStudy (kc_kennylau):

Tell me if there's any step you don't understand

OpenStudy (anonymous):

thx